### v1

# Expert-Guided Prompting and Retrieval-Augmented Generation for Emergency Medical Service Question Answering

Abstract
Large language models (LLMs) have shown promise in medical question answering, yet they often overlook the domain-specific expertise that professionals depend on—such as the clinical subject areas (e.g., trauma, airway) and the certification level (e.g., EMT, Paramedic). Existing approaches typically apply general-purpose prompting or retrieval strategies without leveraging this structured context, limiting performance in high-stakes settings. We address this gap with EMSQA, an 24.3K-question multiple-choice dataset spanning 10 clinical subject areas and 4 certification levels, accompanied by curated, subject area-aligned knowledge bases (40K documents and 2M tokens). Building on EMSQA, we introduce (i) Expert-CoT, a prompting strategy that conditions chain-of-thought (CoT) reasoning on specific clinical subject area and certification level, and (ii) ExpertRAG, a retrieval-augmented generation pipeline that grounds responses in subject area-aligned documents and real-world patient data. Experiments on 4 LLMs show that Expert-CoT improves up to 2.05% over vanilla CoT prompting. Additionally, combining Expert-CoT with ExpertRAG yields up to a 4.59% accuracy gain over standard RAG baselines. Notably, the 32B expertise-augmented LLMs pass all the computer-adaptive EMS certification simulation exams.
Code & Datahttps://uva-dsa.github.io/EMSQA
Extended versionhttps://arxiv.org/pdf/
Introduction
The rapid advancement of large language models (LLMs) has brought new possibilities to high-stakes domains such as emergency medical services (EMS) (Weerasinghe et al. 2024), where accurate and reliable decision-making is critical. There is growing interest in leveraging LLMs for medical education (Abd-Alrazaq et al. 2023), decision support (Ge et al. 2024), and certification preparation (Kung et al. 2023), particularly in the context of open-domain multiple-choice question answering (MCQA). However, while LLMs have shown promising performance on general medical QA benchmarks (Cai et al. 2024), important gaps remain between their current capabilities and the reasoning processes used by trained medical professionals.
Recent approaches in medical MCQA, such as chain-of-thought prompting (CoT) (Wei et al. 2022) and retrieval-augmented generation (RAG) (Lewis et al. 2020), have improved LLM performance by enhancing reasoning capabilities and incorporating external domain knowledge during inference. But these approaches often treat both reasoning and retrieval as undifferentiated processes: the model sees a question and retrieves documents or reasons directly, without considering what kind of knowledge is relevant or how a human expert would approach the task. In contrast, real-world medical professionals typically begin by identifying the subject area of the question—e.g., whether it pertains to trauma, airway management, or pharmacology—and then reason from that domain-specific perspective, using knowledge and protocols appropriate to their level of certification.
Existing benchmarks such as MedQA (Jin et al. 2021), MMLU-Med (Hendrycks et al. 2020) and MedMCQA (Pal, Umapathi, and Sankarasubbu 2022) lack both this structured representation of question expertise (e.g., subject area, certification level) and the associated domain-specific knowledge. This makes it difficult to align retrieval and reasoning processes with human-like problem-solving strategies. In particular, publicly available EMS question answering datasets and knowledge sources with expertise annotation are scarce. Further, current state-of-the-art (SOTA) methods do not account for how incorporating structured expertise can improve the overall effectiveness of reasoning and answer generation in retrieval-augmented systems.
Figure 1: Overall Approach. 1) EMSQA, KB, PR Construction; 2) Tasks: Benchmark LLMs, RAG and Certification Exams.
To address these gaps, we propose a new dataset and a domain expertise-guided LLM framework that models medical MCQA in a more structured, cognitively informed manner. As shown in Figure 1, our contributions are three-fold:
  • We introduce EMSQA, the first EMS MCQA dataset of 24.3K questions, curated based on public and private sources, covering 10 subject areas and 4 certification levels, and accompanied by a structured, subject area-aligned EMS knowledge base (KB) with 40K documents and 4M real-world patient care reports. Partial data (from public sources) and the whole EMS KB is shared as a resource with the EMS and research communities.
  • We propose two approaches to inject domain expertise into LLMs: 1) an expertise-guided prompting strategy (Expert-CoT) that encourages step-by-step reasoning from a domain-specific perspective. 2) an expertise-guided RAG method (ExpertRAG) that selectively retrieves expertise-aligned knowledge from curated EMS KBs and patient records.
  • We benchmark multiple LLMs on EMSQA, evaluating performance across certification levels and subject areas, and compare our framework against SOTA RAG methods. Experimental results show that combining Expert-CoT and ExpertRAG yields up to a 4.59% improvement in accuracy. Notably, the 32B expertise‑augmented models pass all the EMS certification simulation exams.
Related Works
Medical Question Answering Datasets
Medical QA datasets typically fall into two paradigms: retrieval-based tasks (Pampari et al. 2018; Krithara et al. 2023; Ben Abacha and Demner-Fushman 2019), which require explicit evidence grounding by locating answers within documents, and open-domain multiple-choice tasks such as MedQA (Jin et al. 2021), MMLU-Med (Hendrycks et al. 2021), MedMCQA (Pal, Umapathi, and Sankarasubbu 2022) test implicit medical reasoning by choosing the best answer based on world knowledge. However, existing MCQA benchmarks focus on a single certification level and provide subject area labels without corresponding knowledge bases. EMSQA is the first open-domain medical MCQA dataset that spans multiple certification levels while also furnishing both subject area annotations and a structured EMS knowledge base. Table 1 shows a comparison between EMSQA and SOTA open-domain MCQA datasets.
MedQA MedMCQA EMSQA
Domain General Med. General Med. EMS
Data Size 12.7K 193K 24.3K
Exam USMLE AIIMS&NEET PG NREMT
#Certification 1 1 4
#Subject Area 21 10
KB Raw Categorized
Table 1: Comparison of English medical MCQA datasets
Retrieval Augmentation Generation
The basic RAG framework (Lewis et al. 2020) couples a seq2seq model with a dense Wikipedia retriever, and has since been improved via query rewriting (Chan et al. 2024), entity graphs (Edge et al. 2024), and document-structure–aware retrieval (Li et al. 2025), though these methods mainly target general-domain text. For the medical domain, MedRAG (Xiong et al. 2024a) proposes a RAG pipeline with hybrid sparse–dense retrieval on MedCorps, and i-MedRAG (Xiong et al. 2024b) extends it with follow-up question generation and interactive reasoning. Self-BioRAG (Jeong et al. 2024) adapts Self-RAG (Asai et al. 2023) with domain-specific retrieval triggers, while RAG2 (Sohn et al. 2024) leverages rationale-based queries and filtering. EXPRAG (Ou et al. 2025) retrieves similar patient cases, and ClinicalRAG (Lu, Zhao, and Wang 2024) exploits medical entities to query corpora. However, existing medical RAG systems largely ignore question-specific expertise (e.g., subject area or certification level) as an explicit signal to guide retrieval and reasoning.
EMSQA
Data Collection and Preprocessing
We collected a total of 24.3K multiple-choice questions and their corresponding answer choices from practice tests available on 17 websites targeting the National Registry of Emergency Medical Technicians (NREMT) examination (NREMT 2001–2025). The NREMT exam is administered as a Computer Adaptive Test (CAT), which dynamically adjusts question difficulty based on the examinee’s performance, providing a personalized and efficient assessment. It certifies providers at 4 ascending levels of difficulty: entry‐level, Emergency Medical Responder (EMR); basic‐level, Emergency Medical Technician (EMT); intermediate‐level, Advanced EMT (AEMT); and advanced‐level, Paramedic. Our dataset covers all four certification levels. These practice tests assess examinees’ ability to apply medical knowledge, concepts, and principles, as well as their capacity to demonstrate fundamental patient-centered skills.
The overall EMSQA dataset comprises questions from both public and private (subscription-based) websites, but we only release the portion derived from public materials because of copyright restriction of private websites (See Appendix A.2 in Extended version for more details). To ensure the data quality, we did both automatic preprocessing and manual verification of the raw data as shown in Figure 1:
  • Heuristic rules were used to extract and remove special tokens like HTML tags, special symbols.
  • Each question is well-structured and represented as a dictionary containing the following fields: the question text, answer options, correct answer, explanation, source URL, certification level, and subject area.
  • Questions related to images, table were excluded. All the questions are answerable using textual inputs only.
  • All duplicate questions were removed by computing the Levenshtein distance between each pair of questions in the dataset. Pairs with a similarity score greater than 0.9 were considered duplicates and subsequently removed.
  • All questions are manually labeled with their corresponding subject areas, and both questions and answer choices have undergone human proofreading to correct any grammatical errors. A sample subset of 100 questions and KB documents were verified by an EMT expert (See Appendix Section A.5).
Data statistics
As shown in Table 2, the dataset includes a total of 18,602 and 5,669 practice questions from public and private sources, respectively. We use the private questions exclusively for testing and split the public questions into train, validation, and test sets of 13,021, 1,860, 3,721 questions, with average token lengths of 18.27, 19.12, 18.99, respectively. More detailed statistics on preprocessed data are summarized in Appendix Section A.2.2.
Set Size Type Criteria vs KB vs PR
Public Train(13,021) Semantic Avg Sim 79.21 66.45
Val(1,860)
Syntactic
(hit rate)
Vocab 82.95 21.14
Test(3,721) Cpt w/o norm 41.65 8.87
Cpt w/ norm 63.30 15.28
Private Test(5,669) Semantic Avg Sim 80.75 75.35
Syntactic
(hit rate)
Vocab 90.89 28.26
Cpt w/o norm 53.18 14.36
Cpt w/ norm 72.49 22.66
Table 2: Statistics by split for Public and Private EMSQA and Semantic and syntactic evaluation of QA overlap vs. KB/PR. Cpt: Concept; norm: medical normalization.
Figure 2 details the distribution of questions by certification level and subject area. Because certification level is used for evaluation, all questions whose certification level marked with “NA” are relegated to the training split, leaving the validation and test splits to include only questions with explicit EMS certification levels. Subject areas including “airway”, “cardiology”, “EMS operations”, “medical&OB”, and “trauma” dominate the corpus, which mirrors the five domains mandated by the NREMT examination guidelines. The remaining subject areas—“anatomy”, “assessment”, “pharmacology”, “pediatrics”, and “others”—are present as well, but they appear less frequently, reflecting their secondary coverage in the practice materials we collected.
Figure 2: Distributions of Questions by Certification Level (Left) and Subject Area (Right).
Figure 3: Expertise-Guided LLM Framework: Filter training, Expert-CoT Inference, and ExpertRAG Inference.
Knowledge Collection and Preprocessing
To build an external KB for EMSQA, we curated 16 open-access EMS education resources from reputable websites. As shown in Figure 1, these resources span four media types: Youtube video transcripts (Carrie Davis 2025; The EMS Professor 2025), official EMS guidelines (ODEMSA 2025), EMS education textbooks (American Red Cross 2025) or lecture slides (Jones & Bartlett Learning 2025), and EMS flashcards (EMT-Prep 2025). For Youtube transcripts, we segment each transcript into sections using the title cues supplied by the uploader. The PDF documents were converted to plain text with PyPDF2 and each file was parsed chapter-by-chapter using page ranges gleaned from its table of contents. We then leveraged GPT-4o (Achiam et al. 2023) to reorganize the raw chapter text into a coherent section-level hierarchy (See the prompt in Appendix Section A.3.1). Lastly, we manually audited the output by checking that every section heading and its span of text aligned with the corresponding passage in the original PDF and corrected any discrepancies. Based on the chapter name, we manually split our collected EMS KB into 10 subject areas including “airway&ventilation”, “anatomy”, “assessment”, “cardiovascular”, “ems operations”, “medical&ob”, “pediatrics”, “pharmacology”, “trauma” and “others”. After cleaning, our KB comprises 39,652 documents, contains 2,545,192 tokens with 34,110 vocabularies.
To evaluate the helpfulness of the KB for EMSQA, we assess our KB’s coverage from both syntactic and semantic aspects. For semantic evaluation, we leverage MedCPT (Jin et al. 2023) to retrieve, for each question, its most similar document from the KB and report the average similarity score between each question and its retrieved document. For syntactic evaluation, we removed stop words and computed vocabulary hit rate (KB $\cap$ QA / QA) between the KB and EMSQA. We also follow the methods in  (Ge et al. 2024) to extract EMS-related concepts defined in (Preum et al. 2020), both before and after UMLS normalization (Bodenreider 2004). Table 2 reports both syntactic and semantic overlaps with our KB. Syntactic hit rates span from 41.65% to 90.89%. Semantic similarity, measured by average cosine similarity, was 79.21% on public and 80.75% on private data. This indicates that our KB can support answering most EMSQA questions. Appendix Section A.3 provides comprehensive details on web crawling, EMS concept extraction, KB statistics, and the overlap between EMSQA and the KB.
Patient Care Report Collection and Preprocesing
We used the National EMS Information System (NEMSIS) (Dawson 2006) 2021 public research dataset as our source of patient records (PR). NEMSIS is a large tabular corpus in which each record contains field information such as dispatch details, scene information, initial assessment, EMS protocols and triage, vital signs, medications and procedural interventions, and patient history. As shown in Figure 1, we removed any field whose value was “NA,” “Not Applicable,” “Not Recorded,” or “Unknown.” If more than 30% of a record’s fields were discarded, we excluded the entire record. Finally, we converted each remaining record into plain text by concatenating its key–value pairs. After cleaning, our corpus comprised 4,003,430 records with an average token length of 311.7. Based on the protocol field, we further categorized the patient records into 6 subject areas including “airway”, “assessment”, “cardiovascular”, “medical&ob”, “pediatrics”, and “trauma”.
As shown in Table 2, we evaluate NEMSIS coverage of EMSQA data along two dimensions: syntactic hit‐rate and semantic similarity. Semantic coverage remains high, with patient records achieving average similarity scores of 66.45% on the public split and 75.35% on the private split of EMSQA. In contrast, syntactic concept hit‐rate range from 8.87% to 28.26%, substantially lower than those observed for the KB. Detailed fields extracted from NEMSIS and their summary statistics are presented in Appendix Section A.4.
Methodology
Task Formulation
Given the $i$th question $q_{i}$, its answer options $\mathcal{O}_{i}=\{o_{1},\dots,o_{m}\}$, along with the expertise annotations of subject area $s_{i}$ and certification level $l_{i}$, the goal of MCQA is to maximize the likelihood of selecting the correct answer $a_{i}^{*}\in\mathcal{O}_{i}$. Let $\mathcal{R}$ be a retriever function that takes $q_{i}$ as input and returns a set of relevant documents. A language model $f$ selects an answer $A_{i}$ as:
$A_{i}=\arg\max_{o\in\mathcal{O}_{i}}f(o\mid q_{i},\mathcal{O}_{i},\mathcal{R}(q_{i}))$ (1)
We propose an expertise-guided LLM framework (see Figure 3) with two strategies for incorporating domain expertise in this task: (1) Expert-CoT, a prompting strategy that encodes $s_{i}$ and $l_{i}$ into a prompt template to guide LLM reasoning based on question-specific expertise; and (2) ExpertRAG, a RAG pipeline that retrieves expert-aligned knowledge conditioned on $s_{i}$. During inference, we use a trained Filter to predict subject area $\hat{s}_{i}$ and certification level $\hat{l}_{i}$ for prompt encoding and knowledge retrieval.
Filter Training
To guide the LLM reasoning and RAG retrieval based on question-specific expertise, we train a lightweight LLM-based filter to predict the key expertise attributes, including question’s subject area and certification level. As shown in Figure 3 (Left), we adopt LoRA (Hu et al. 2022) to inject a small set of trainable parameters into the model while keeping the full LLM weights fixed. We augment the LoRA modules with two classification heads, $W_{\mathrm{sub}}$ and $W_{\mathrm{lvl}}$, that predict the question’s subject area and certification level, and optimize them jointly in a multi‐task setting. We append a special token <classify> at the end of each query, and extract the hidden state $h_{i}$ of this final token from the last layer and feed it into our two classification heads:
$h_{i}$ $=\mathrm{LM}_{\mathrm{last}}(q_{i},\mathcal{O}_{i}\|\langle\mathrm{classify}\rangle),$ (2)
$(p_{i}^{\mathrm{sub}},\,p_{i}^{\mathrm{lvl}})$ $=\big(\sigma(W_{\mathrm{sub}}^{\!\top}h_{i}),\,\sigma(W_{\mathrm{lvl}}^{\!\top}h_{i})\big)$
where $\sigma$ is sigmoid function. We use binary cross‐entropy for subject area classification and cross‐entropy for certification classification. We set hyper-parameters $w_{\mathrm{sub}}$ and $w_{\mathrm{lvl}}$ to balance the multi-task loss. The overall training objective is
$\mathcal{L}=w_{\mathrm{sub}}\cdot\mathrm{BCE}\left({p}_{i}^{\mathrm{sub}},{y}_{i}^{\mathrm{sub}}\right)+w_{\mathrm{lvl}}\cdot\mathrm{CE}\left(p_{i}^{\mathrm{lvl}},y_{i}^{\mathrm{lvl}}\right)$ (3)
During inference, given a question $q_{i}$ with answer options $\mathcal{O}_{i}$, the filter first predicts a certification-level probability distribution $p_{i}^{\mathrm{lvl}}$ and a multi-label subject area probability vector $p_{i}^{\mathrm{sub}}$. The predicted certification level and subject area set are computed as:
$\hat{s}_{i}=\mathbf{1}\{p_{i}^{\mathrm{sub}}>0.5\},\quad\hat{l}_{i}=\arg\max p_{i}^{\mathrm{lvl}}.$ (4)
Expertise-Guided Prompting (Expert-CoT)
Figure 3 (Middle) illustrates our Expert-CoT prompting method. Standard CoT prompts encourage LLMs to reason step by step but do not specify where to begin. In contrast, Expert-CoT prompting guides the model’s reasoning by explicitly providing the subject area and certification level as starting point for the thought process. The final answer is generated by passing the predicted subject area $\hat{s}_{i}$ and certification level $\hat{l}_{i}$, into the Expert-CoT prompt template:
$\hat{A}_{i}=f^{\mathrm{CoT\text{-}Expert}}\bigl(q_{i},\;\mathcal{O}_{i},\;\hat{l}_{i},\;\hat{s}_{i}\bigr).$ (5)
Expertise-Guided RAG (ExpertRAG)
As shown in Figure 3 (Right), for ExpertRAG the filter’s predicted subject area guides the retriever to search for relevant knowledge base entries and patient records tailored to the question’s subject area. The LLM then conditions on the predicted expertise and the retrieved documents to generate the final answer. Based on $q_{i}$ and the predicted subject area $\hat{s}_{i}$, we explore three retrieval strategies for retriever $\mathcal{R}$:
  • Global: Retrieve the top $M$ and $N$ evidence documents from the entire KB and PR, respectively. This serves as a baseline corresponding to standard RAG without any subject area filtering.
  • Filter then Retrieve (FTR): First filter the whole KB and PR to retain only documents matching the predicted subject area $\hat{s}_{i}$, then retrieve the top $M$ and $N$ documents from these filtered subsets.
  • Retrieve then Filter (RTF): First retrieve a larger candidate set from the whole KB and PR (e.g., $10\times M$ from KB and $10\times N$ from PR), then filter out documents whose subject area do not match $\hat{s}_{i}$, retaining the top $M$ and $N$ relevant documents.
The final answer is generated by passing the retrieved documents, along with the predicted subject area and certification level, into the RAG prompt template:
$\hat{A}_{i}=f^{\mathrm{RAG}}\bigl(q_{i},\;\mathcal{O}_{i},\;\mathcal{R}(q_{i},\hat{s}_{i}),\;\hat{l}_{i},\;\hat{s}_{i}\bigr).$ (6)
Experiments
We conduct extensive experiments to evaluate Expert-CoT and ExpertRAG methods by applying them to different baseline LLMs and comparing their performance to SOTA LLMs and RAGs. We aim to answer three research questions:
RQ1: Where do SOTA LLMs shine or stumble on EMSQA across subject areas and certification levels?
RQ2: How much does explicit expertise injected by Expert-CoT and ExpertRAG lift baseline accuracy?
RQ3: Can expertise-aware LLMs meet the NREMT passing score at different certification levels?
LLM Baselines
We use three categories of SOTA baseline models: 1) Open-source LLMs, we select Qwen3-32B (Team 2025), and LLama-3.3-70B (Grattafiori et al. 2024) because of their great performance in multiple domains; 2) Medical LLMs, we select OpenBioLLM-70B (Ankit Pal 2024), which currently leads the Open Medical-LLM Leaderboard (Pal et al. 2024). 3) Closed-source LLMs, we choose OpenAI-o3 (Brown et al. 2020) and Gemini-2.5-pro (Team et al. 2023), both of which achieve top results on a range of benchmarks. We further apply 0 to 64-shot and CoT prompting  (Wei et al. 2022) and Expert-CoT to each baseline to benchmark the performance under different prompt strategies.
RAG Baselines
We select the following SOTA medical RAG models due to their superior performance and code availability: 1) MedRAG (Xiong et al. 2024a), which is a RAG toolkit combining multiple medical documents; 2) i-MedRAG (Xiong et al. 2024b), which iteratively refines medical queries via multi-step retrieval; 3) Self-BioRAG (Jeong et al. 2024), which self-reflectively decides when to retrieve biomedical texts and then generates the answer; 4) Qwen3-4B + KB, which is a vanilla RAG pipeline with our collected KB as retrieval corpora; 5) Qwen3-4B + PR, a vanilla RAG pipeline with our collected PR as retrieval corpora; 6) Qwen3-4B + Global, a vanilla RAG pipeline with both KB and PR as retrieval corpora. We also include 7) Qwen3-4B with 0-shot and 8) Qwen3-4B with CoT prompting as baselines. For a fair comparison, we applied RAG with Qwen3-4B across all methods, except Self-BioRAG, which is trained from scratch. All baseline RAG methods used CoT prompting.
Implementation Details
For ExpertRAG, we use Qwen3-4B as the core LLM. We employ MedCPT (Jin et al. 2023) as our retriever due to its superior performance in the medical domain and its widespread use in SOTA RAG systems. We fix the number of retrieved documents to $M=32$ for KB retrieval and $N=8$ for PR retrieval. We chunk all KB and PR documents using a fixed window size of 512 tokens with an overlap of 128 tokens. Since some questions in our MCQA dataset have multiple correct answers, we report both exact-match accuracy (Acc) and sample-based F1 over all answer options (F1) (Khashabi et al. 2018; Wang et al. 2019).
To train the filter, we fine-tune LoRA modules with rank r=8r{=}8, scaling factor $\alpha{=}16$, and a dropout rate of 0.05, using the sequence length of 128 tokens. To balance the certification and subject area classification objectives, we apply DWA (Liu, Johns, and Davison 2019) with $T=2$ to dynamically adjust $w_{\mathrm{cat}}$ and $w_{\mathrm{lvl}}$. We use AdamW (Loshchilov and Hutter 2017) optimizer and regularization with a weight decay of 0.01. We set the decision threshold of 0.5 for subject area classification. We fix the random seed at 42, and run all experiments on NVIDIA H200 GPUs.
Experimental Results
Filter Performance
Since subject area classification is a multi-label task and certification level classification is a multi-class task, we evaluate the classification by micro f1-score (miF) and macro f1-score (maF). The performance of filter and other baselines (0-shot, 4-shot, and CoT) are shown at Table 3. Results show our filter trained with LoRA with two classification heads significantly outperforms the baselines.
Split Model Method Subject Area Certification
miF maF miF maF
Public Filter LoRA 80.72 71.92 65.87 63.45
Qwen3-4B 0-shot 55.43 51.61 45.77 30.49
Qwen3-4B 4-shot 56.33 54.42 45.46 30.28
Qwen3-4B CoT 59.72 55.66 47.80 35.77
Private Filter LoRA 79.06 70.48 65.54 63.50
Qwen3-4B 0-shot 42.93 31.73 44.12 25.01
Qwen3-4B 4-shot 45.76 34.08 44.04 29.41
Qwen3-4B CoT 46.22 35.49 47.70 31.92
Table 3: Filter’s Performance on Expertise Classification
LLM Benchmarking and Expert-CoT Evaluation
To evaluate the strengths and limitations of SOTA LLMs on EMSQA (RQ1) and to assess how domain expertise, injected via Expert-CoT, influences reasoning (RQ2), we benchmark multiple LLMs under different prompting strategies. Table 4 and Figure 4 present the results. We highlight several key findings from this evaluation:
Figure 4: (a) Certification level 0-shot performance. (b) Subject area 0-shot performance on the Public dataset
Model Prompt Public Private
Acc F1 Acc F1
OpenBioLLM 0-shot 57.67 57.76 63.86 64.76
CoT 59.88 60.34 67.01 67.77
 (GT) Expert-CoT 61.92 62.03 68.75 69.82
 (Filter) Expert-CoT 61.32 61.93 67.79 68.32
Llama-3.3 0-shot 81.69 82.69 78.06 78.77
CoT 81.89 83.08 85.16 86.35
 (GT) Expert-CoT 82.42 83.35 86.49 87.62
 (Filter) Expert-CoT 82.40 83.18 86.63 87.65
Qwen3-32B 0-shot 83.55 83.55 85.11 85.89
4-shot 84.41 84.41 85.48 86.13
32-shot 81.13 81.13 82.22 83.41
64-shot 82.48 82.48 86.22 87.26
CoT 84.96 84.97 88.78 90.13
 (GT) Expert-CoT 85.70 85.71 89.73 90.98
 (Filter) Expert-CoT 85.57 85.60 89.50 91.20
OpenAI-o3 0-shot 92.39 92.39
Gemini-2.5 0-shot 89.36 89.36
Table 4: Accuracy and F1 (%) of LLMs under Public vs. Private Data. GT/Filter: Ground-truth/predicted expertise.
Model Description Public Private
Acc F1 Acc F1
No-RAG Baselines
Qwen3-4B 0-shot 70.99 71.01 69.88 69.95
Qwen3-4B CoT 72.35 73.09 70.58 72.02
RAG Baselines + CoT
MedRAG RAG on Med 74.31 74.41 71.12 73.33
i-MedRAG Iterative RAG 77.96 78.00 74.02 76.35
Self-BioRAG SelfRAG on Bio 55.71 58.84 45.72 49.67
Qwen3-4B KB 76.49 76.07 75.02 76.53
Qwen3-4B PR 73.02 73.96 70.54 72.38
Qwen3-4B Global 78.12 79.17 75.46 76.87
RAG Baselines + Expert-CoT $\Delta_{\mathrm{Acc/F1}}=\mathord{+}1.38/\mathord{+}0.46$
Qwen3-4B KB 78.02 79.04 76.01 76.25
Qwen3-4B PR 73.82 73.82 71.53 72.96
Qwen3-4B Global 79.59 79.61 76.75 77.35
ExpertRAG-GT + CoT $\Delta_{\mathrm{Acc/F1}}=\mathord{+}3.35/\mathord{+}2.71$
ExpertRAG FTR 80.97 81.34 79.13 80.00
ExpertRAG RTF 81.11 81.45 79.17 80.01
ExpertRAG-GT + Expert-CoT $\Delta_{\mathrm{Acc/F1}}=\mathord{+}4.59/\mathord{+}3.69$
ExpertRAG FTR 81.62 81.65 80.40 81.02
ExpertRAG RTF 82.24 82.26 80.51 81.16
ExpertRAG-Filter + Expert-CoT $\Delta_{\mathrm{Acc/F1}}=\mathord{+}3.44/\mathord{+}2.59$
ExpertRAG FTR 80.99 80.99 79.45 80.16
ExpertRAG RTF 80.95 80.96 79.47 80.22
Table 5: End-to-end RAG Performance and Ablation Study on ExpertRAG and Expert-CoT: “GT” = Ground-truth expertise; “Filter” = Predicted expertise.
Model Description EMR EMT AEMT Paramedic
Pass Score Acc T Pass Score Acc T Pass Score Acc T Pass Score Acc T
Qwen3-4B 0-shot 809 64.18 33 940 74.07 33 940 71.64 33 940 71.72 35
Expert-CoT 940 72.42 59 940 76.73 73 1179 80.41 76 940 76.03 87
ExpertRAG-4B FTR+Expert-CoT 1218 84.21 66 940 78.76 61 940 77.31 69 940 79.67 74
RTF+Expert-CoT 940 76.47 59 1185 81.30 93 1190 83.53 67 940 77.61 86
Qwen3-32B 0-shot 1207 82.65 22 1140 81.63 23 1280 85.92 26 1163 81.58 29
Expert-CoT 1261 86.27 50 1255 86.96 52 1310 89.11 61 1292 89.01 57
ExpertRAG-32B FTR+Expert-CoT 1350 92.22 75 1292 89.01 76 1215 84.60 86 1228 83.93 125
RTF+Expert-CoT 1350 92.22 75 1328 92.32 82 1356 92.31 82 1276 88.04 99
Table 6: Pass (✓) or Fail (✗) Summary of Models by Simulation Certification Test. T: Overall Time (min).
Closed-source models outperform open-source models. In particular, OpenAI-o3 consistently achieves the highest overall accuracy of 92.39. Among open-source models, Qwen3-32B achieves the best accuracy of 85.70, though a significant gap remains compared to closed-source models.
Few-shot prompting improves accuracy up to a point. We varied the number of in‑context exemplars from 0 to 64 (See Appendix Section A.7.2) and observed that incorporating a small number of examples yields substantial gain over the zero‑shot baseline. However, adding examples beyond a certain point leads to diminishing or no improvement.
Expert-CoT help guide LLM reasoning. Integrating domain expert knowledge via CoT‑Expert guides reasoning towards appropriate context and consistently boosts CoT prompting performance by up to 2.05% across models. Also, using the Filter’s predicted expertise for Expert-CoT leads to comparable performance to when using ground-truth.
Baseline LLMs underperform on easier questions. As shown in Figure 4a, across all models, we observe that performance is lowest for the EMR certification level (the most basic tier in the NREMT exam) and highest for the Paramedic level. This may be due to smaller data size for EMR level and the procedural nature of EMR questions, whereas Paramedic questions aligning more closely with the academic-style content seen during LLM pretraining.
LLMs excel in pharmacology and anatomy but falter on core NREMT domains. Figure 4b shows that models reliably answer pharmacology and anatomy questions but perform poorly on trauma, operations, and medical&OB. One reason may be task complexity. The questions in the former subject areas need single-hop, fact-based queries which can be easily solved in a zero-shot manner, whereas the latter demand multi-hop reasoning and richer EMS context.
Ablation Study on ExpertRAG and Expert-CoT
To study how much does expertise help LLMs (RQ2), we report an ablation over six configurations at Table 5: (1) No-RAG, (2) RAG+CoT with a standard global retriever on EMS PR and KB, (3) RAG+Expert-CoT, (4) ExpertRAG-GT+CoT, (5) ExpertRAG-GT+Expert-CoT, and (6) ExpertRAG-Filter+Expert-CoT. The best configuration outperforms the baseline by 4.59 / 3.69 points in Acc / F1.
Effect of PR and KB. Row 2 vs. row 1 isolates the gain of adding PR and KB as retrieval documents with standard global retriever. Results show KB brings more improvement than PR, and combing both yields the best performance.
Effect of Expert-CoT. Row 3 vs. row 2 ablates the additional gain from Expert-CoT, showing that expertise-aware reasoning is better than standard reasoning.
Effect of Expert-RAG. Rows 4 vs. 2 (and 5 vs. 3) ablate the impact of our Expert-RAG (FTR/RTF) compared with standard RAG with a global retriever. With ground-truth subject area and certification level, ExpertRAG consistently outperforms SOTA RAG baselines, highlighting the value of expertise-guided retrievers. Across settings, both FTR and RTF outperform global retrieval, and RTF achieves the best overall performance.
Effect of Filter. Row 6 vs. 5 measures the effect of using Filter’s predicted expertise rather than ground-truth expertise: although there is a small performance drop, ExpertRAG-Filter still surpasses the best baseline.
NREMT Computer Adaptive Simulation Tests
To investigate whether our best models can be certified in NREMT exam (RQ3), we subscribed to MedicTests (MedicTests 2025), the NREMT Computer Adaptive Simulation Test. The simulation exam consists of 80-150 questions with adaptive difficulty and needs to be completed in 2.5 hours. The NREMT cognitive exam uses a scaled score to determine passing status. The passing score is 950 on a scale of 100 to 1500. We evaluated our Expert-CoT and ExpertRAG models and 0-shot baselines on the NREMT simulation exam. The expertise-augmented models use the trained filter to predict the subject area and certification. The Pass or Fail result is shown at Table 6. All models complete the exam within the allotted time with expertise-augmented models taking much longer times. Expertise-augmented models consistently lead to higher test scores and significantly improve accuracy compared to baseline LLMs. However, performance varies with model size. Specifically, 4B models fail at one or more certification levels, whereas larger 32B models pass all four. Among these, ExpertRAG‑32B with the RTF retrieval strategy achieves the highest overall score across certification levels. Importantly, although smaller models do not meet the certification threshold, their scores (940) closely approach the passing bar (950).
Conclusion
This paper proposes a domain expertise-aware LLM framework for medical multiple-choice question answering that incorporates expertise to guide both LLM reasoning and RAG retrieval. We introduce EMSQA, the first large-scale labeled MCQA dataset for EMS with subject area and certification-level annotations, and curated EMS knowledge bases. We propose Expert-CoT, which guides LLM reasoning from a domain-informed perspective, and ExpertRAG, which retrieves expertise-specific knowledge for augmented generation. Experiments show that integrating expertise into LLM reasoning and RAG retrieval significantly improves performance over baselines. More importantly, our expertise-augmented models pass the NREMT simulation tests for all EMS certification levels. EMSQA provides a new benchmark for MCQA research in medical domain and our proposed expertise-augmented LLM framework can be applied to other medical MCQA datasets with similar or other expertise attributes.
Ethics Statement
Although our 32B expertise-augmented model is able to pass the simulated NREMT examination, all models studied in this work are research prototypes and are not approved medical devices. They must not be used as the sole basis for diagnosis or treatment decisions for real patients. Any outputs should only be used as a reference by licensed healthcare professionals, who remain fully responsible for clinical judgment and patient care. The models may generate incorrect, incomplete, or biased recommendations and may not reflect up-to-date guidelines. Our experiments are conducted entirely in simulation, and no decisions produced by the models were used to influence real-world patient care.
Acknowledgments
This work was supported by Award 70NANB21H029 from the U.S. Department of Commerce, National Institute of Standards and Technology (NIST).
Ethics Statement
Although our 32B expertise-augmented model is able to pass the simulated NREMT examination, all models studied in this work are research prototypes and are not approved medical devices. They must not be used as the sole basis for diagnosis or treatment decisions for real patients. Any outputs should only be used as a reference by licensed healthcare professionals, who remain fully responsible for clinical judgment and patient care. The models may generate incorrect, incomplete, or biased recommendations and may not reflect up-to-date guidelines. Our experiments are conducted entirely in simulation, and no decisions produced by the models were used to influence real-world patient care.
Acknowledgments
This work was supported by Award 70NANB21H029 from the U.S. Department of Commerce, National Institute of Standards and Technology (NIST).
Appendix
A.1 Introduction
This is the technical appendix for the paper ”Expert-Guided Prompting and Retrieval-Augmented Generation for Emergency Medical Service Question Answering”.
Source # QAs Certification Level(s) Explanation
Public
(Bluefield Rescue 2025) 955 EMT
(NREMT Practice Test 2025) 240 EMT
(SmartMedic 2025) 537 EMT
(Union Test Prep 2025) 150 EMT
(Montgomery County MD 2025) 130 EMT
(Mometrix Academy 2025) 138 EMT, AEMT, Paramedic
(PracticeTestGeeks 2025) 99 EMR
(My CPR Certification Online 2025) 74 EMR, EMT, AEMT
(Quizizz 2025) 1,554 EMR, EMT, AEMT, Paramedic
(MinniQuiz 2025) 4,444 EMR, EMT, Paramedic
(Quizlet 2025) 9,084 EMR, EMT, AEMT, Paramedic, NA
(CareerEmployer 2025) 100 AEMT
(DocDrop / Arthur Hsieh 2025) 456 EMT
(LearningExpress Hub 2025) 1,336 Paramedic, EMT
(PocketPrep 2025) 47 EMR, EMT, AEMT, Paramedic
Private
(EMT-Prep 2025) 4,843 EMR, EMT, AEMT, Paramedic, Critical Care
(Jones & Bartlett Learning 2025) 1,285 EMT
Table 7: Summary of NREMT practice question sources in EMSQA.
A.2 EMSQA Data
A.2.1 EMSQA Data Collection
In Tables 7, we present detailed information on 17 public and private EMSQA resources, including each source’s name and URL, number of questions, certification level and availability of explanations.
A.2.2 EMSQA Data Statistics
In Table 8, we show the detailed EMSQA dataset statistics. The dataset includes a total of 18,602 and 5,669 practice questions from public and private sources, respectively. We use the private questions exclusively for testing and split the public questions into train, validation, and test sets of 13,021, 1,860, 3,721 questions, with average token lengths of 18.27, 19.12, 18.99, respectively.
Data Split
#Explanations
#Choices
(avg/max)
#Answers
(avg/max)
Question Tokens
(avg/max)
Choice Tokens
(avg/max)
Tokens
Vocab
Public Train (13,021) 2217 4.01 / 7.00 1.00 / 3.00 18.27 / 218 6.28 / 240 565,303 14,017
Val (1,860) 383 3.99 / 5.00 1.00 / 3.00 19.12 / 155 6.01 / 44 80,215 6,629
Test (3,721) 773 4.01 / 6.00 1.00 / 3.00 18.99 / 135 6.10 / 60 161,464 8,913
Total (18,602) 3132 4.01 / 7.00 1.00 / 3.00 18.50 / 218 6.22 / 240 806,982 16,032
Private Test (5,669) 5451 4.01 / 6.00 1.06 / 4.00 30.44 / 355 5.46 / 47 296,673 10,637
Table 8: Statistics by split for Public and Private EMSQA
A.3 Knowledge Base Preprocessing & Statistics
PROMPT:
Well formated the unstructured text by the folllowing rules:
1. Fix awkward or broken line breaks within sentences or paragraphs. 2. Separate paragraphs with a single blank line. 3. Remove figure captions and references (e.g., “Figure 8–1”, “see Figure…”). 4. Remove page numbers (e.g., “239”). 5. Remove attribution or copyright lines, such as: defined copyrights 6. **Do not rephrase, reword, rewrite, or summarize; do not add any other words in your response**. 7. If the unstructured text contains subtitles, treat each subtitle as a key and store its corresponding paragraph(s) as the value. Return only the resulting JSON. Here is the unstructured text to format: RAW TEXT
Figure 5: Prompt for organizing chapter text
A.3.1 Prompt for Organizing Chapter Text
As shown in Figure 5, we present the prompt that instructs the GPT-4o to divide each chapter’s raw text into clearly labeled sections.
A.3.2 Prompt for EMS Concept Extraction
In main paper Table 3: Semantic and syntactic evaluation of QA overlap vs. KB/PR, we present the concept overlap in Syntactic metrics, we present the prompt we used for EMS concept extraction in Figure 6.
PROMPT:
Extract all EMS concepts from the following text.
Here are some examples of EMS concepts: gelastic epilepsy, visual seizure, pallor, pale color, aox4, pare down, cut down Return all the extracted EMS concepts as a lowercase letter in strict JSON format, like: [”fever”, ”cardiac arrest”] Here is one example: Text: Early symptoms include cough, wheezing, shortness of breath. Lung transplantation is an option. Response: Let’s think step by step, Step1: label the tokens one by one ”EMS concept”, or ”none”. -Early: none -symptoms: none -include: none -cough: EMS concept -wheezing: EMS concept -shortness: EMS concept -of: none -breath: EMS concept -Lung: EMS concept -transplantation: EMS concept -is: none -an: none -option: none Step2: RefineEMS concept from Step 1 by following criteria, 1.concatenate EMS concept spans 2.remove extra irrelevant words in EMS concept -cough: EMS concept -wheezing: EMS concept -shortness of breath: EMS concept -Lung transplantation: EMS concept Step3: Return the your result: [”cough”, wheezing”, ”shortness of breath”, ”Lung transplantation”] Now is the real text: RAW TEXT
Figure 6: Prompt for EMS Concept Extraction
Data Comparison QA KB PR KB $\cap$ QA PR $\cap$ QA KB $\setminus$ QA PR $\setminus$ QA KB $\cup$ QA PR $\cup$ QA
Public Vocab 15,892 33,965 6,773 13,183 3,359 20,782 3,414 36,674 19,306
Cpts (w/o norm) 25,594 61,003 9,383 10,661 2,269 50,342 7,114 75,936 32,708
Cpts (w norm) 17,262 34,627 8,088 10,927 2,637 23,700 5,451 40,962 22,713
Private Vocab 10,496 33,965 6,773 9,540 2,966 24,425 3,807 34,921 14,303
Cpts (w/o norm) 11,621 61,003 9,383 6,180 1,669 54,823 7,714 66,444 19,335
Cpts (w norm) 8,689 34,627 8,088 6,299 1,969 28,328 6,119 37,017 14,808
Table 9: Detailed overlap statistics between QA and KB for public and private data.
Resource Title Level # Vocab
Transcripts
Emergency Care and Transportation of
the Sick and Injured Advantage Package (Carrie Davis 2025)
EMT 23,836
Nancy Caroline’s Emergency Care in
the Streets (Carrie Davis 2025)
Paramedic
AAOS Advanced Emergency Medical
Technician (AEMT) 4th Ed (The EMS Professor 2025)
AEMT
AAOS Critical Care Transport Paramedic (The EMS Professor 2025) Paramedic
Textbooks
EMT Exam for Dummies (DocDrop / Arthur Hsieh 2025) EMT 2,025
Emergency Medical Services: Clinical
Practice and Systems Oversight (Wiley 2025)
18,799
EMS Essentials: A Resident’s Guide
to Prehospital Care(EMRA 2025)
4,123
Emergency Medical Response - RedCross (American Red Cross 2025) EMR 10,096
Emergency Medical Responder: Your First
Response in Emergency Care (Firearms Training Los Angeles CA 2025)
EMR 7,794
Guidelines
Professional Responder Cheat Sheet (MediPro First Aid 2025) EMR 1,569
EMS Pharmacology Reference Guide (Rhode Island Department of Health 2025) 3,072
Paramedic Medication Manual (Delaware Health and Social Services 2025) Paramedic 2,011
ODEMSA Regional EMS Documents (ODEMSA 2025) 6,340
Slides
Emergency Care and Transportation of
the Sick and Injured Advantage Package (Jones & Bartlett Learning 2025)
EMT 8,597
Flashcards & Study Guides
EMS Study Guides (EMT-Prep 2025) EMR, EMT, AEMT, Paramedic 8,373
Table 10: Knowledge base collection details
A.3.3 UMLS Concept Normalization
We applied UMLS api (Bodenreider 2004) to normalize the extracted EMS concepts and retained only terms whose semantic types fell into one of the following categories: Sign or Symptom (Sosy), Finding (Fndg), Laboratory or Test Result (Lbtr), Clinical Attribute (Clna), Quantitative Concept (Qnco), Qualitative Concept (Qlco), Disease or Syndrome (Dsyn), Mental or Behavioral Dysfunction (Mobd), Pathologic Function (Patf), Neoplastic Process (Neop), Congenital Abnormality (Cgab), Anatomical Abnormality (Anab), Injury or Poisoning (Inpo), Cell or Molecular Dysfunction (Celf), Body Part/Organ/Organ Component (Bpoc), Body Location or Region (Bodl), Body Space or Junction (Bsoj), Body System (Bodsys), Therapeutic or Preventive Procedure (Topp), Diagnostic Procedure (Diap), Laboratory Procedure (Lbpr), Imaging Procedure (Impr), Health Care Activity (Hlca), Clinical Drug (Clnd), Pharmacologic Substance (Phsu), Antibiotic (Antb), Vitamin (Vita), Organic Chemical (Orch), Amino Acid/Peptide/Protein (Aapp), Biologically Active Substance (Bacs), Hazardous or Poisonous Substance (Hops), Steroid (Strd), Hormone (Horm), Medical Device (Medd), Temporal Concept (Tmco), Spatial Concept (Spco), and Functional Concept (Fngp).
A.3.4 Knowledge and EMSQA Overlap Statistics
The detailed statistics such as the total number of concepts in EMSQA and KB, overlapped concepts (KB $\cap$ QA), union concepts (KB $\cup$ QA), KB unique concepts (KB $\setminus$ QA) are shown in Table 9. The syntactic (hit rate) reported in paper is calculated by KB $\cap$ QA / QA.
A.3.5 Knowledge Base Collection Details
Table 10 shows the detailed information of every knowledge source we collected to construct external EMS knowledge bases. Specifically, we list the url, source type, source title, certification level, number of tokens and vocabulary of each source. In total, there are 2,545,192 tokens in our external knowledge bases, and the vocabulary size 34,110.
A.4 NEMSIS Patient Care Report Preprocessing & Statistics
A.4.1 NEMSIS Patient Record Preprocessing
To concatenate information for each patient records, we used the following fields in different tables, and concatenate by the shared primary key. The keys in our patient records include ”Date - Time of Symptom Onset”, ”Chief Complaints”, ”Possible Injury”, ”Cause of Injury”, ”Chief Complaint Anatomic Location”, ”Chief Complaint Organ System”, ”Gender”, ”Age”, ”Age Unit”, ”Level of Responsiveness (AVPU)”, ”Primary Symptoms”, ”Other Associated Symptoms”, ”Primary Impressions”, ”Secondary Impressions”, ”Protocol Age Category”, ”Protocols”, ”Date - Time Vital Signs Taken”, ”ECG Type”, ”SBP (Systolic Blood Pressure)”, ”Heart Rate”, ”Respiratory Rate”, ”Pulse Oximetry”, ”Blood Glucose Level”, ”End Tidal Carbon Dioxide (ETCO2)”, ”Glasgow Coma Score-Eye”, ”Glasgow Coma Score-Verbal”, ”Glasgow Coma Score-Motor”, ”Pain Scale Score”, ”Stroke Scale Score”, ”Stroke Scale Type”, ”Reperfusion Checklist”, ”Date - Time Medication Administered”, ”Medication Administered Prior to this Unit’s EMS Care”, ”Medication Given”, ”Medication Dosage”, ”Medication Dosage Units”, ”Response to Medication”, ”Role - Type of Person Administering Medication”, ”Date - Time Procedure Performed”, ”Procedure”, ”Number of Procedure Attempts”, ”Response to Procedure”, ”Role - Type of Person Performing the Procedure”, ”Alcohol - Drug Use Indicators”, ”Barriers to Patient Care”, ”Cardiac Arrest”, ”Date - Time of Cardiac Arrest”, ”First Monitored Arrest Rhythm of the Patient”, ”Cardiac Arrest Etiology”, ”Type of CPR Provided”, ”Cardiac Rhythm on Arrival at Destination”, ”Reason CPR - Resuscitation Discontinued”, ”Incident - Patient Disposition”, ”EMS Transport Method”, ”Transport Mode from Scene”, ”Initial Patient Acuity”, ”Final Patient Acuity”.
We further classify patient records to seven subject areas by the field protocol EMS. Specifically,
”Assessment” includes ”General-Universal Patient Care/ Initial Patient Contact”, ”General-Individualized Patient Protocol”
Medical & OB includes ”Medical-Seizure”, ”Medical-Nausea/Vomiting”, ”Medical-Influenza-Like Illness/ Upper Respiratory Infection”, ”Medical-Abdominal Pain”, ”Medical-Altered Mental Status”, ”OB/GYN-Pregnancy Related Emergencies”, ”Medical-Supraventricular Tachycardia (Including Atrial Fibrillation)”, ”Medical-Cardiac Chest Pain”, ”Medical-Tachycardia”, ”Medical-Syncope”, ”Medical-Stroke/TIA”, ”Medical-Respiratory Distress/Asthma/COPD/Reactive Airway”, ”Medical-Respiratory Distress-Bronchiolitis”, ”Medical-Diarrhea”, ”Medical-Hypoglycemia/Diabetic Emergency”, ”Medical-Hyperglycemia”, ”Medical-Allergic Reaction/Anaphylaxis”, ”Medical-Hypertension”, ”Medical-Bradycardia”, ”Medical-Pulmonary Edema/CHF”, ”Medical-Hypotension/Shock (Non-Trauma)”, ”Medical-Opioid Poisoning/Overdose”, ”Medical-ST-Elevation Myocardial Infarction (STEMI)”, ”Medical-Ventricular Tachycardia (With Pulse)”, ”Medical-Stimulant Poisoning/Overdose”, ”Medical-Respiratory Distress-Croup”, ”Medical-Adrenal Insufficiency”, ”Medical-Beta Blocker Poisoning/Overdose”, ”Medical-Calcium Channel Blocker Poisoning/Overdose”, ”OB/GYN-Gynecologic Emergencies”, ”OB/GYN-Childbirth/Labor/Delivery”, ”OB/GYN-Eclampsia”, ”OB/GYN-Post-partum Hemorrhage”, ”General-Overdose/Poisoning/Toxic Ingestion”, ”General-Fever”, ”General-Epistaxis”, ”General-Back Pain”, ”General-Pain Control”, ”Environmental-Altitude Sickness”, ”Environmental-Cold Exposure”, ”Environmental-Hypothermia”, ”Environmental-Heat Stroke/Hyperthermia”, ”Environmental-Heat Exposure/Exhaustion”
”Airway” includes ”Airway”, ”Airway-Failed”, ”Airway-Sedation Assisted (Non-Paralytic)”, ”Airway-Rapid Sequence Induction (RSI-Paralytic)”, ”Airway-Obstruction/Foreign Body”,
”EMS Operations” includes ”Exposure-Airway/Inhalation Irritants”, ”General-Neglect or Abuse Suspected”, ”Exposure-Explosive/ Blast Injury”, ”General-IV Access”, ”General-Refusal of Care”, ”General-Behavioral/Patient Restraint”, ”General-Interfacility Transfers”, ”General-Spinal Immobilization/Clearance”, ”General-Medical Device Malfunction”, ”General-Law Enforcement - Assist with Law Enforcement Activity”, ”General-Extended Care Guidelines”, ”General-Exception Protocol”, ”General-Indwelling Medical Devices/Equipment”, ”General-Community Paramedicine / Mobile Integrated Healthcare”, ”General-Law Enforcement - Blood for Legal Purposes”, ”General-Dental Problems”, ”Exposure-Biological/Infectious”, ”Exposure-Carbon Monoxide””, ”Exposure-Chemicals to Eye”, ”Exposure-Smoke Inhalation”, ”Exposure-Cyanide”, ”Exposure-Radiologic Agents”, ”Exposure-Blistering Agents”, ”Exposure-Nerve Agents”
”Cardiovascular” includes ”General-Cardiac Arrest”, ”Cardiac Arrest-Asystole”, ”Cardiac Arrest-Determination of Death / Withholding Resuscitative Efforts”, ”Cardiac Arrest-Pulseless Electrical Activity”, ”Cardiac Arrest-Do Not Resuscitate”, ”Cardiac Arrest-Ventricular Fibrillation/ Pulseless Ventricular Tachycardia”, ”Cardiac Arrest-Post Resuscitation Care”, ”Cardiac Arrest-Special Resuscitation Orders”, ”Cardiac Arrest-Hypothermia-Therapeutic”
”Pediatrics” includes ”Medical-Apparent Life Threatening Event (ALTE)”, ”Medical-Newborn/ Neonatal Resuscitation”
”Trauma” includes ”Environmental-Frostbite/Cold Injury”, ”Injury-Head”, ”Injury-Extremity”, ”Injury-Burns-Thermal”, ”Injury-General Trauma Management”, ”Injury-Multisystem”, ”Injury-Eye”, ”Injury-Mass/Multiple Casualties”, ”Injury-Amputation”, ”Injury-Spinal Cord”, ”Injury-Conducted Electrical Weapon (e.g., Taser)”, ”Injury-Bleeding/ Hemorrhage Control”, ”Injury-Bites and Envenomations-Land”, ”Injury-Facial Trauma”, ”Injury-Cardiac Arrest”, ”Injury-Crush Syndrome”, ”Injury-Thoracic”, ”Injury-Drowning/Near Drowning”, ”Injury-Diving Emergencies”, ”Injury-Electrical Injuries”, ”Injury-Topical Chemical Burn”, ”Injury-Impaled Object”, ”Injury-Bites and Envenomations-Marine”, ”Injury-Lightning/Lightning Strike”, ”Injury-SCUBA Injury/Accidents”.
NEMSIS Patient Records
Total cases 4,003,430
Total tokens 1,247,811,207
Average tokens/case 311.7
Min tokens 202
Max tokens 824
Vocabulary 6,838
Table 11: Statistics for the NEMSIS patient care reports
A.4.2 NEMSIS Patient Record Statistics
The detailed statistics of processed NEMSIS data is shown at Table 11. There are in total 4,003,430 patient records with average 311.7 tokens per case. The total number vocabulary is 6,838.
We also show the statistics such as the total number of concepts in EMSQA and PR, overlapped concepts (PR $\cap$ QA), union concepts (PR $\cup$ QA), PR unique concepts (PR $\setminus$ QA) are shown in Table 9. The syntactic (hit rate) reported in paper is calculated by PR $\cap$ QA / QA.
A.5 Human Evaluation of EMSQA and Knowledge Base
To validate the quality and clinical utility of our EMSQA benchmark and its accompanying curated knowledge bases, we conducted an expert review on a randomly sampled subset of 100 questions. Sampling was stratified to ensure balanced coverage across certification levels—25 questions each for EMR, EMT, AEMT, and Paramedic—and across five major medical subject areas—20 questions each for Airway, Trauma, Cardiology, Medical, and Operations—while also drawing uniformly from all source links.
We asked one EMT-Certified professionals rated each question–answer pair on four dimensions:
Clinical Accuracy
Verify that the provided “correct” answer is clinically sound. Rating: Yes / No
Difficulty Level
Judge whether the question’s difficulty matches the expected level for its certification. Rating: 1 (Strongly Disagree)–5 (Strongly Agree)
Subject Area Accuracy
Confirm that the assigned medical subject area appropriately reflects the question’s content. Rating: 1 (Strongly Disagree)–5 (Strongly Agree)
Knowledge Relevance
Assess whether the assembled knowledge base sufficiently supports answering the question. Rating: 1 (Strongly Disagree)–5 (Strongly Agree)
Table 12 shows average rating for four dimensions. The results indicate that: (1) the provided answers are clinically accurate; (2) the EMT expert largely agrees with the assigned certification levels and subject areas; and (3) the curated knowledge base is helpful for answering the questions.
Dimension Scale Expert Rating
Clinical Accuracy 0-100 100
Difficulty Level 1-5 3.95
Subject Area Accuracy 1-5 4.85
Knowledge Relevance 1-5 4.39
Table 12: Human Expert evaluation on EMSQA & KB.
Relevance
Strategy Hit@k P@k MRR MAP
(1 / 5 / 10) (1 / 5 / 10)
Global 44.0 / 69.0 / 78.0 44.0 / 35.0 / 31.7 55.6 39.4
FTR 44.0 / 71.0 / 80.0 44.0 / 35.0 / 32.2 57.6 39.6
RTF 45.0 / 75.0 / 81.0 45.0 / 35.0 / 32.5 57.5 39.6
Supportiveness
Strategy Hit@k P@k MRR MAP
(1 / 5 / 10) (1 / 5 / 10)
Global 25.0 / 52.0 / 67.0 25.0 / 21.0 / 20.3 38.3 27.6
FTR 27.0 / 59.0 / 70.0 27.0 / 24.0 / 21.2 39.7 29.2
RTF 29.0 / 58.0 / 70.0 29.0 / 23.0 / 21.6 39.8 28.8
Table 13: Retrieval Quality Across Relevance and Supportiveness
EMSQA—Public
Model Prompt EMR EMT AEMT Paramedic Overall
Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
Qwen3-32B 0-shot 79.08 79.08 82.98 82.98 83.68 83.68 86.69 86.69 83.55 83.55
4-shot 80.89 80.89 84.31 84.31 86.05 86.05 86.06 86.06 84.41 84.41
8-shot 80.20 80.20 84.60 84.60 85.79 85.79 86.14 86.14 84.39 84.39
16-shot 81.03 81.03 84.23 84.23 85.79 85.79 87.31 87.31 84.82 84.82
32-shot 77.68 77.68 81.13 81.13 81.58 81.58 83.01 83.01 81.13 81.13
64-shot 80.33 80.33 80.09 80.09 85.26 85.26 85.43 85.43 82.48 82.48
CoT 80.06 80.06 82.90 82.95 85.26 85.26 86.92 86.92 84.96 84.97
 (GT) Expert-CoT 82.43 82.48 85.49 85.49 87.89 87.89 87.16 87.16 85.70 85.71
 (Filter) Expert-CoT 81.52 81.51 82.24 83.68 86.37 86.65 86.98 87.51 85.07 85.10
Llama-3.3 0-shot 79.08 79.08 81.05 81.05 85.53 85.53 85.67 85.67 81.69 82.69
CoT 79.64 81.29 79.57 80.79 83.42 84.51 85.20 86.13 81.89 83.08
 (GT) Expert-CoT 79.22 80.32 80.68 81.77 83.68 85.15 85.75 86.24 82.42 83.35
 (Filter) Expert-CoT 80.20 81.24 80.24 81.16 82.63 83.51 85.83 86.27 82.40 83.18
OpenBioLLM 0-shot 51.60 51.73 54.18 54.26 61.58 61.71 63.66 63.74 57.67 57.76
CoT 53.02 53.03 54.78 54.96 62.34 62.56 63.81 63.97 59.88 60.34
 (GT) Expert-CoT 55.31 55.89 57.82 57.87 64.66 65.03 65.77 65.81 61.92 62.03
 (Filter) Expert-CoT 54.91 54.91 55.12 55.12 65.16 65.17 65.52 65.54 61.32 61.93
OpenAI-o3 0-shot 90.79 90.79 93.49 93.49 92.37 92.37 92.17 92.17 92.39 92.39
Gemini-2.5 0-shot 87.59 87.59 89.93 89.93 90.26 90.26 89.51 89.51 89.36 89.36
EMSQA—Private
Qwen3-32B 0-shot 84.81 84.93 84.55 85.28 85.55 86.47 85.27 86.09 85.11 85.89
4-shot 86.60 86.72 85.41 85.95 86.60 87.30 85.54 86.24 85.48 86.13
8-shot 85.26 85.39 84.62 85.20 85.91 86.64 85.71 86.42 85.39 86.07
16-shot 85.78 85.95 84.98 85.55 86.11 86.84 85.68 86.41 85.52 86.19
32-shot 82.58 82.95 81.80 82.84 83.69 85.12 82.92 84.26 82.22 83.41
64-shot 87.71 87.90 85.75 86.74 87.52 88.89 87.13 88.19 86.22 87.26
CoT 91.81 92.01 87.86 89.08 90.11 91.66 89.83 91.24 88.78 90.13
 (GT) Expert-CoT 94.42 94.74 88.63 89.85 92.01 93.58 91.59 92.83 89.73 90.98
 (Filter) Expert-CoT 91.66 92.97 88.42 90.05 91.60 93.08 91.11 92.60 89.50 91.20
Llama-3.3 0-shot 75.95 76.12 77.93 78.63 76.87 77.66 76.75 77.47 78.06 78.77
CoT 87.57 88.16 84.65 85.86 86.88 88.04 85.93 87.00 85.16 86.35
 (GT) Expert-CoT 89.72 90.48 85.90 87.04 88.66 89.61 87.55 88.58 86.49 87.62
 (Filter) Expert-CoT 90.17 90.70 86.06 86.98 88.41 89.47 87.64 88.63 86.63 87.65
OpenBioLLM 0-shot 60.98 61.24 60.96 61.74 66.25 67.23 67.32 68.31 63.86 64.76
CoT 62.16 62.73 64.35 64.89 69.85 70.71 72.13 72.42 67.01 67.77
 (GT) Expert-CoT 63.60 63.83 65.01 65.88 70.87 71.84 73.69 74.74 68.75 69.82
 (Filter) Expert-CoT 62.26 62.43 64.31 64.78 70.01 70.78 72.01 72.23 67.79 68.32
Table 14: Zero-/few-shot, CoT and Expert-CoT results for LLMs on EMSQA.
A.6  Retrieval Strategy Performance
In the paper we present three different retrieval strategies (Global, Filter then Retrieve (FTR), Retrieve then Filter (RTF)), to evaluate the retrieval performance, we uniformly sampled 100 questions from different 4 certification levels (25 questions per certification level) and 10 subject areas ($\sim$ 10 questions per subject area). Then we use MedCPT to retrieve 32 documents from Knowledge Bases.
Since we don’t have the ground-truth question-document pair, we use LLM-as-a-judge (Zheng et al. 2023) to evaluate the Relevance (Es et al. 2024; Asai et al. 2023) and Supportiveness (Asai et al. 2023) of retrieved 32 documents for every question. Relevance judges if each retrieved chunk is relevant to the question. And supportiveness judges if this retrieved chunk support to answer the question.
PROMPT:
Question: {}
Answer: {} Retrieved Chunk: {} Rubric - Relevance • ”Relevant” — The chunk provides information that is useful for answering the question. • ”Irrelevant” — The chunk is off-topic or supplies no helpful information. - Supportiveness Evaluate how much of a hypothetical answer could be *entailed* by this chunk alone. • ”Fully” — All necessary information is present; the answer would be completely supported. • ”Partially” — Some but not all required information is present. • ”None” — No information is supported, or the chunk contradicts the answer.
Figure 7: Prompt for LLM-as-a-judge
Figure 8: Zero-shot Performance per Subject Area on EMSQA-Private Dataset
As shown in Figure 7, we present the prompt template we used in LLM-as-a-judge. To evaluate “Relevance” and “Supportiveness” we use four Information Retrieval metrics: Top-$k$ Hit rate ($\mathit{hit}@k$), which measures the fraction of questions for which at least one positive chunk appears in the top-$k$ retrieved; Top-$k$ Precision ($p@k$), which computes the average proportion of positive chunks among the top-$k$ for each question; Mean Reciprocal Rank ($\mathrm{MRR}$), which takes the reciprocal of the rank position of the first positive chunk for each question and then averages these reciprocals—thereby reflecting how early in the ranking the first positive appears; and Mean Average Precision ($\mathrm{MAP}$), which for each question computes the average of precisions at every positive chunk’s position and then averages those per‐question APs, capturing both the ranking quality and distribution of positives over the entire list. We omit Top-$k$ Recall ($r@k$) since we lack ground-truth documents.
For each question $q$ in sampled questions $\mathcal{Q}$ and retrieved chunk $i$, define
$r_{q,i}^{\mathrm{rel}}=\begin{cases}1&\text{if chunk $i$ is Relevant},\\ 0&\text{otherwise},\end{cases}$ (7)
$r_{q,i}^{\mathrm{sup}}=\begin{cases}1&\text{if chunk $i$ is Fully or Partially support},\\ 0&\text{otherwise}.\end{cases}$ (8)
Then, applied to either $\{r_{q,i}^{\mathrm{rel}}\}$ or $\{r_{q,i}^{\mathrm{sup}}\}$:
$\mathit{hit}@k=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\mathbf{1}\!\Bigl(\sum_{i=1}^{k}r_{q,i}>0\Bigr)$ (9)
$p@k=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\frac{1}{k}\sum_{i=1}^{k}r_{q,i}$ (10)
$\mathrm{MRR}=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\frac{1}{\min\{\,i:r_{q,i}=1\}}$ (11)
$\mathrm{MAP}=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\frac{1}{\sum_{i=1}^{N}r_{q,i}}\sum_{i=1}^{N}r_{q,i}\,\frac{1}{i}\sum_{j=1}^{i}r_{q,j}\,.$ (12)
Table 13 presents the full retrieval performance of three strategies—Global (direct retrieval), Retrieve-then-Filter (RTF) and Filter-then-Retrieve (FTR)—on both “Relevance” and “Supportiveness.” In the Relevance dimension, RTF and FTR consistently outperform Global across all metrics, showing that proposed retrieval strategies retrieve relevant chunks earlier and more reliably. Likewise, in the Supportiveness dimension, both RTF and FTR achieve higher Hit@k, P@k, MRR and MAP than Global, indicating they better capture fully or partially supportive evidence. Overall, these results demonstrate that our proposed retrieval strategies more effectively extract both relevant and supportive context from the knowledge base, which in turn provides stronger grounding for the LLM’s final answer.
A7. Benchmark LLMs
A.7.1 Zero-shot Performance per Subject Area
As shown in Figure 8, we show the zero-shot LLMs’ per subject area performance on private dataset. We only run open-source models on our private dataset. The conclusion remains consistent on the private dataset: LLMs excel in pharmacology and anatomy but falter on core NREMT domains. Models answer pharmacology and anatomy items reliably, yet stumble on trauma, airway management, EMS operations, and medical&OB questions. One reason may be task complexity: the former are largely single-hop, fact-based queries, whereas the latter demand multi-hop reasoning and richer EMS context.
A.7.2 Detailed statistics of Benchmarks
Table 14 summarizes the full benchmarking results for each model and prompting strategy, reporting accuracy and F1 across all four EMS certification levels (EMR, EMT, AEMT, Paramedic) as well as overall performance. We list several key findings:
Closed-source models outperform open-source models. In particular, OpenAI-o3 consistently achieves the highest overall accuracy of 92.39. Among open-source models, Qwen3-32B achieves the best accuracy of 85.70, though a significant gap remains compared to closed-source models.
Few-shot prompting improves accuracy up to a point. We varied the number of in‑context exemplars from 0 to 64 (See Appendix Section A.7.2) and observed that incorporating a small number of examples yields substantial gain over the zero‑shot baseline. However, adding examples beyond a certain point leads to diminishing or no improvement.
Expert-CoT help guide LLM reasoning. Integrating domain expert knowledge via CoT‑Expert guides reasoning towards appropriate context and consistently boosts CoT prompting performance by up to 2.05% across models. Also, using the Filter’s predicted expertise for Expert-CoT leads to comparable performance to when using ground-truth.
Model Description Public Private
Acc F1 Acc F1
Qwen3-32B CoT 84.96 84.97 88.78 90.13
+ Subject Area 85.36 85.48 88.94 90.61
+ Certification 85.40 85.45 88.89 90.62
Expert-CoT 85.70 85.71 89.73 90.98
Qwen3-4B RAG 78.12 79.17 75.46 76.87
+ Subject Area 81.11 81.45 79.17 80.07
+ Certification 79.16 79.21 76.38 77.25
ExpertRAG 82.24 82.26 80.51 81.16
Table 15: Ablation Study: Effect of Expertise (Certification, Subject Area) on Expert-CoT and ExpertRAG
A8. Ablation Studies on Certification and Subject Area
To assess which expertise signal contributes more to performance, we ablate two signals—Subject Area and Certification—and measure their effects on both Expert-CoT and ExpertRAG. The results are presented in Table 15.
Both expertise help, and combining them is best.. Both “Subject Area” and “Certification” help to improve performance in Expert-CoT and ExpertRAG, and combing both yields the best performance.
“Subject Area” matters more in RAG. “Subject Area” has more improvement than “Certification” in RAG. In our retrieval pipelines (RTF/FTR), “Subject Area” is used to filter irrelevant KB and PR documents, directly improving retrieved context quality. “Certification” is applied only in prompting, so it has no effect on retrieval and thus yields smaller performance gains in ExpertRAG.
A9. Error Analysis
Figure 9: Error Analysis
We conducted an error analysis on our best model, ExpertRAG-GT + Expert-CoT, which achieved 82.24% accuracy. From the 661 errors on the test set, we randomly selected 50 samples to manually examine the questions, retrieved knowledge, and model reasoning. As shown in Figure 9, we attribute the errors to five main categories:
1) Reasoning Error (42.4%): the retrieved documents are helpful to answer the question and the model used the knowledge but the reasoning is incorrect.
2) Retrieval Error(20.3%): no relevant knowledge is retrieved, and model reasoning by “model knowledge””, leading to the wrong answer.
3) Fail to Use the Correct Knowledge (20.3%): part of the retrieved knowledge is helpful to answer the question but the model use other irrelevant knowledge for reasoning.
4) Misinterpret the Question (15.3%): the model didn’t interpret the question correctly. For example, one question is what is the age cutoff for using an aed on a patient? and the choices are a. 18 year old, b. 5 year old, c. 25 year old, d. 1 year old. The question is asking about minimum age (d. 1 year old) for routine AED use. However the model’s understanding is adult age threshold (a. 18 year old).
5) Model Policy Constraint (1.7%): the model think none of the choices is correct and abstain to answer.